OpenStudy (anonymous):

How to solve (dy/dx)-(e^x) (sinx) =0? Separation of variables and integration.

OpenStudy (experimentx):

just solve it ... dy/dx = e^x sin(x) dy = e^x sin(x) dx integrate both sides

OpenStudy (anonymous):

Wait hold on srry that exponent of e is "y" not "x"

OpenStudy (anonymous):

I integrated (1/e^y)dy=sinxdx, but then...is that -u^2/2=-ln(cosx)+c???

OpenStudy (anonymous):

and so e^-2y=2ln|cosx|+c

OpenStudy (anonymous):

and then I came to e^-2y=ln(cosx^2) +c and I don't really know how to deal with this...

hartnn (hartnn):

\(\large \int \dfrac{1}{e^y}dy=\int e^{-y}dy = ....?\) try this integral again ?

hartnn (hartnn):

thats left side integral for the right side, its simply \(\Large \int \sin x dx\)

OpenStudy (anonymous):

and the integral of sinxdx is -ln(cosx)+c

hartnn (hartnn):

\(\Large \int \sin x dx = -\cos x +c\)

OpenStudy (anonymous):

ohhh maybe I was thinking of tangetn hahah uhm... and the e^-y, use u -substitution?

hartnn (hartnn):

i am not sure if you're allowed to use this directly, \(\large \int \dfrac{1}{e^y}dy=\int e^{-y}dy = \dfrac{e^{-y}}{-1}+c\) dividing by the co-efficient of the variable

OpenStudy (anonymous):

???

hartnn (hartnn):

you didn't get that ? then use a u-substitution as -y = u

OpenStudy (anonymous):

ohhhhhhh!

OpenStudy (anonymous):

so then -e^-y=cosx+c, then ln both sides?

hartnn (hartnn):

you need to isolate 'y' ? then yes!

OpenStudy (anonymous):

uhm... where does the negative on the e go when we put the ln on?

hartnn (hartnn):

first isolate e^(-y) (note that right side is actually -cos x + c and not just cos x +c) so, e^(-y) = cos x - c now take ln on both sides

OpenStudy (anonymous):

oh! right! I completely forgot the negative on the cosine... so then y=ln(cosx)^-1 +c?

hartnn (hartnn):

lets see \(\large \ln (e^{-y}) = \ln (\cos x -c)\)

hartnn (hartnn):

\(\large -y \ln e = \ln (\cos x -c)\)

hartnn (hartnn):

and since ln e = 1

hartnn (hartnn):

\(\large y = -\ln (\cos x -c )\) thats it!

OpenStudy (anonymous):

yay! okay cool~ ^^ thank you!

hartnn (hartnn):

welcome ^_^